ELECTRONIC TRANSITIONS OF TUNGSTEN MONOXIDE

\underline{GANG LI}, TSUYOSHI HIRAO, and PETER F. BERNATH, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

Electronic transitions of tungsten monoxide were detected in emission and several new bands were discovered in the region 13,000 - 24,500 cm^{−1}using a Bruker IFS 120 HR high resolution Fourier transform spectrometer. Results of rotational analyses of the 0-0 bands of various systems and related discussions are presented.